Draw the shear and moment diagrams. 20 kN/m 80 kN/m A D B C I 3 m 1 m 1 m
Answer #1

20KNIM 160 kN/m A 120 I B с 3 Support Teactons RatRD20 (4) + < (3) (60) : 140 EMA ZO 2014)(2) +2 (3) (60)(1+3,0)) - R(5) zo 4Moment 2 Bending @AB @BC Mr = 84x - 202 Mx = 843.203*(20)(2-1) (2-1) at x=1 Mg:74 KN- al xoo MAC at xel Me: 865N-M at er ] MB